OpenStudy (anonymous):

Which expression is the simplified form of the square root of 49a^4 7a 7a^2 7a * the square root of a 7a^2 * the square root of a

OpenStudy (anonymous):

\[\large \sqrt{49a^4} = \sqrt{7 \times 7 \times a^2 \times a^2} =7a^2\]

OpenStudy (anonymous):

okay thats what i thought, just checking. thank you!

OpenStudy (anonymous):

It is okay.. Have faith on yourself..